OpenStudy (anonymous):

This is my last problem:) Find the time t in years when the annual sales x of a new product are increasing at the greatest rate. Use a graphing utility to verify your results. Please help:) Will give medal

OpenStudy (anonymous):

\[x=\frac{ 600000t^2 }{ 36+ t^2 }\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you in calculus?

OpenStudy (anonymous):

Not sure, but I can't reply to messages =/ I'm in Business Calc

OpenStudy (anonymous):

I'm not sure how to start the problem

jimthompson5910 (jim_thompson5910):

are you able to find the derivative of that function?

OpenStudy (anonymous):

I believe is (36+t^2)(1200000)-(2T)(T^2

OpenStudy (anonymous):

I may be completely wrong

jimthompson5910 (jim_thompson5910):

close, (36+t^2)(1,200,000t) - (600000t^2)(2t) all over (36+t^2)^2

jimthompson5910 (jim_thompson5910):

this function represents how much x is changing with respect to t we want to find the greatest rate of change, so we have to derive again because we want to locate the local max of this function

jimthompson5910 (jim_thompson5910):

to help make things a bit simpler, the derivative simplifies to (43200000t)/(36+t^2)^2

jimthompson5910 (jim_thompson5910):

now you have to derive (43200000t)/(36+t^2)^2

OpenStudy (anonymous):

this part I believe is where I got stuck

jimthompson5910 (jim_thompson5910):

what do you get when you derive (43200000t)/(36+t^2)^2 ?

OpenStudy (anonymous):

(43200000)/(36+t^2)^4 ?

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

Let u = 43,200,000t v = (36+t^2)^2 so, du/dt = ?? dv/dt = ??

OpenStudy (anonymous):

du/dt = 43200000 dv/dt = I'm not sure about this one

jimthompson5910 (jim_thompson5910):

use the chain rule to find dv/dt

OpenStudy (anonymous):

I git 144t + 3t^2

jimthompson5910 (jim_thompson5910):

v = (36+t^2)^2 dv/dt = 2(36+t^2)*(2t) dv/dt = 4t(36+t^2) dv/dt = 4t^3+144t

jimthompson5910 (jim_thompson5910):

use du/dt and dv/dt to find the derivative of (43200000t)/(36+t^2)^2

OpenStudy (anonymous):

I honestly don't understand

OpenStudy (anonymous):

none of the problem except the first part

OpenStudy (anonymous):

my Professor never went over this problem at all

jimthompson5910 (jim_thompson5910):

so you've never gone over local min, local max and things like that?

OpenStudy (anonymous):

not deep into like now we just when over basic problems, not word

jimthompson5910 (jim_thompson5910):

The idea of this entire problem is this We have some sales function x(t). The derivative, x ' (t), represents how fast sales occur instantaneously at any given time t. This derivative function is an upside down parabola. It peaks out at some point. This peak is the point we're interested in. This is the point where the speed peaks out and reaches its maximum. How do we find this peak? We derive x ' (t) to get x '' (t) and set that equal to 0. We want to find when the tangent has a horizontal line.

jimthompson5910 (jim_thompson5910):

So we have x ' (t) = (43200000t)/(36+t^2)^2 derive it using the quotient rule u = 43200000t u' = 43200000 v = (36+t^2)^2 v' = 4t^3+144t

jimthompson5910 (jim_thompson5910):

x'(t) = u/v x''(t) = (u'*v - v'*u)/( v^2 ) x''(t) = [ 43200000*(36+t^2)^2 - (4t^3+144t)*(43200000t)]/( (36+t^2)^2)^2 x''(t) =-(129600000*(-12+t^2))/(36+t^2)^3 I skipped a bunch of steps, but this gives you the basic idea

jimthompson5910 (jim_thompson5910):

Now we set x''(t) equal to zero and solve for t -(129600000*(-12+t^2))/(36+t^2)^3 = 0 -129600000*(-12+t^2) = 0 t = ??? or t = ???

OpenStudy (anonymous):

The computer gives me 2sqrt 3

jimthompson5910 (jim_thompson5910):

same here

jimthompson5910 (jim_thompson5910):

so at t = 2*sqrt(3) which is approximately t = 3.464101615, the sales speed is maxed out

jimthompson5910 (jim_thompson5910):

to find what the speed is at t = 3.464101615, you plug t = 3.464101615 into x'(t) which is the derivative function

OpenStudy (anonymous):

Thank you i will write all this down to keep it for future reference:) I have never seen a problem like this before..

jimthompson5910 (jim_thompson5910):

to find the amount of sales done at that point in time, you plug t = 3.464101615 into x(t). You may have to round

OpenStudy (anonymous):

Thank You!

jimthompson5910 (jim_thompson5910):

you're welcome